If you invest $1,000 per period for the following number of periods, how much would you have at the end? a. 10 years at 10 percent. b. 20 years at 9 percent.
Answer #1

a

Future value = present value*(1+ rate)^time
Future value = 1000*(1+0.1)^10
Future value = 2593.74
Using Calculator: press buttons "2ND"+"FV" then assign
PV =-1000
I/Y =10
N =10
PMT = 0
CPT FV
Using Excel
=FV(rate,nper,pmt,pv,type)
=FV(0.1,10,,-1000,)

b

Future value = present value*(1+ rate)^time
Future value = 1000*(1+0.09)^20
Future value = 5604.41
Using Calculator: press buttons "2ND"+"FV" then assign
PV =-1000
I/Y =9
N =20
PMT = 0
CPT FV
Using Excel
=FV(rate,nper,pmt,pv,type)
=FV(0.09,20,,-1000,)